You want to be able to withdraw $35,000 each year for 15 years. Your account earns 6% interest.
a) How much do you need in your account at the beginning?
$

b) How much total money will you pull out of the account?
$

c) How much of that money is interest?
$

Answer by Boreal(15235)

You will pull out of the account $35,000*15=$525,000
-
PV=35000(1-(1.06)^(-15))/0.06, round at end=$339,928.71
It is what you want*(1-(1+(r/n)^nt/(r/n). Here, r=0.06 and n=1
-
the amount that is interest is the amount you pull out minus what you had at the beginning.
That is $185,071.29.